OpenStudy (anonymous):

simplify by factoring √x^16

OpenStudy (saifoo.khan):

x^8

OpenStudy (anonymous):

I typed that in and it says first factor √x^16 *√x

OpenStudy (saifoo.khan):

\[\Huge \sqrt{x^{16}}\]u have that right?

OpenStudy (anonymous):

I got the answer

OpenStudy (anonymous):

x^8√x

OpenStudy (saifoo.khan):

u miswrote the question before. hehe

OpenStudy (saifoo.khan):

\[\sqrt{x^{16}} \times \sqrt {x} \to x^8 \sqrt x\]

OpenStudy (anonymous):

yeah that's what I forgot lol

OpenStudy (anonymous):

I am so awful with algebra

OpenStudy (saifoo.khan):

hehe. Algebra is easy! :O

OpenStudy (anonymous):

lol I wish it was for me! Too many rules lol

OpenStudy (saifoo.khan):

Rofl. Correct. haha

OpenStudy (anonymous):

lol